# Location–Scale Calibration for Generalized Posterior

### v2

@@ Appendix B Proofs of the main results
Write $\varphi(\boldsymbol{x})=\mathcal{N}(\boldsymbol{x}\mid\boldsymbol{0},\boldsymbol{H}_{0}^{-1})$. Using the change of variables $\boldsymbol{y}=\boldsymbol{x}+\boldsymbol{\delta}_{n}$ and the triangle inequality,
@@
$\int_{\mathbb{R}^{p}}\Bigl|q_{n}^{\tilde{\boldsymbol{\theta}}_{n}}(\boldsymbol{x})-\mathcal{N}\!\bigl(\boldsymbol{x}|\boldsymbol{0},\boldsymbol{H}_{0}^{-1}\bigr)\Bigr|\mathop{}\!\mathrm{d}\boldsymbol{x}\longrightarrow_{p}0.$
Recalling that $\boldsymbol{H}_{0}=\eta\boldsymbol{J}_{\lambda}^{\star}$, this is precisely the claimed total-variation convergence of the law of $\sqrt{s_{n}}(\boldsymbol{\theta}-\tilde{\boldsymbol{\theta}}_{n})$ under $\Pi_{n}^{\eta}$ to $\mathcal{N}(\boldsymbol{0},(\eta\boldsymbol{J}_{\lambda}^{\star})^{-1})$ at any admissible center $\tilde{\boldsymbol{\theta}}_{n}$. ∎
@@
$\sqrt{s_{n}}\bigl(\boldsymbol{\theta}^{(d)}-\boldsymbol{\theta}_{\mathrm{GB}}\bigr)\longrightarrow_{d}\mathcal{N}\bigl(\boldsymbol{0},\boldsymbol{H}_{0}^{-1}\bigr),\quad n\to\infty.$ (19)
@@
$\sqrt{s_{n}}\bigl(\boldsymbol{\theta}_{\mathrm{calib}}^{(d)}-\tilde{\boldsymbol{\theta}}_{n}\bigr)=\boldsymbol{\Omega}\sqrt{s_{n}}\bigl(\boldsymbol{\theta}^{(d)}-\boldsymbol{\theta}_{\mathrm{GB}}\bigr),$ (20)
@@
$\sqrt{s_{n}}\bigl(\boldsymbol{\theta}_{\mathrm{calib}}^{(d)}-\tilde{\boldsymbol{\theta}}_{n}\bigr)\longrightarrow_{d}\mathcal{N}\bigl(\boldsymbol{0},\boldsymbol{\Omega}\boldsymbol{H}_{0}^{-1}\boldsymbol{\Omega}^{\top}\bigr),\quad n\to\infty.$ (21)
@@
$\sqrt{s_{n}}\bigl(\boldsymbol{\theta}_{\mathrm{calib}}^{(d)}-\tilde{\boldsymbol{\theta}}_{n}\bigr)\longrightarrow_{d}\mathcal{N}\bigl(\boldsymbol{0},\boldsymbol{V}_{\mathrm{target}}^{\star}\bigr).$
@@ Appendix C Proofs of the plug-in results
Let $q_{n}^{\bar{\boldsymbol{\theta}}_{n}}$ be the density of the scaled posterior $\boldsymbol{X}_{n}=\sqrt{s_{n}}(\boldsymbol{\theta}-\tilde{\boldsymbol{\theta}}_{n})$ under $\Pi_{n}^{\eta}$ for an admissible center $\tilde{\boldsymbol{\theta}}_{n}$. We first present two generic facts that will be used repeatedly.
@@
$\int_{\mathbb{R}^{p}}\left|q_{n}^{\bar{\boldsymbol{\theta}}_{n}}(\boldsymbol{x})-\left.\mathcal{N}\bigl(\boldsymbol{x}\right|\boldsymbol{0},\boldsymbol{H}_{0}^{-1}\bigr)\right|\mathop{}\!\mathrm{d}\boldsymbol{x}\longrightarrow 0,\quad n\to\infty$
and that there exist random constants $C_{n},c_{n}>0$, bounded in probability, such that
$q_{n}^{\bar{\boldsymbol{\theta}}_{n}}(\boldsymbol{x})\leq C_{n}\exp\!\left(-c_{n}|\boldsymbol{x}|^{2}\right)\quad\text{for all }\boldsymbol{x}\in\mathbb{R}^{p},\text{ all }n.$
Then, for any function $g:\mathbb{R}^{p}\to\mathbb{R}^{k}$, ($k\in\mathbb{N}$) with polynomial growth,
$\int g(\boldsymbol{x})q_{n}^{\bar{\boldsymbol{\theta}}_{n}}(\boldsymbol{x})\mathop{}\!\mathrm{d}\boldsymbol{x}\to\int g(\boldsymbol{x})\left.\mathcal{N}\bigl(\boldsymbol{x}\right|\boldsymbol{0},\boldsymbol{H}_{0}^{-1}\bigr)\mathop{}\!\mathrm{d}\boldsymbol{x}.$
@@
$\mathbb{E}\!\left[\boldsymbol{X}_{n}\right]\longrightarrow\boldsymbol{0},\quad\mathbb{E}\!\left[\boldsymbol{X}_{n}\boldsymbol{X}_{n}^{\top}\right]\longrightarrow\boldsymbol{H}_{0}^{-1},\quad\sup_{n}\mathbb{E}\!\left[\|\boldsymbol{X}_{n}\|^{4}\right]<\infty.$
@@
On $\{\|\boldsymbol{x}\|\leq R\}$, $g$ is bounded, so total-variation convergence yields convergence of the first term as $n\to\infty$, and the limit as $R\to\infty$ recovers the Gaussian expectation. For the tail term, polynomial growth of $g$ together with the Gaussian domination implies
$\sup_{n}C_{n}\int_{\|\boldsymbol{x}\|>R}\bigl|g(\boldsymbol{x})\bigr|\exp(-c_{n}\|\boldsymbol{x}\|^{2})\mathop{}\!\mathrm{d}\boldsymbol{x}\longrightarrow 0,\quad R\to\infty.$
[…]
@@
$\sqrt{s_{n}}\bigl(\boldsymbol{\theta}_{\mathrm{GB}}-\tilde{\boldsymbol{\theta}}_{n}\bigr)\longrightarrow_{p}\boldsymbol{0}$
@@
$s_{n}\operatorname{Var}_{\Pi_{n}^{\eta}}\!\left[\boldsymbol{\theta}\right]=\mathbb{E}\!\left[\boldsymbol{X}_{n}\boldsymbol{X}_{n}^{\top}\right]+o_{p}(1),$
where $\boldsymbol{X}_{n}=\sqrt{s_{n}}(\boldsymbol{\theta}-\tilde{\boldsymbol{\theta}}_{n})$.
@@
$\boldsymbol{\theta}_{\mathrm{GB}}=\mathbb{E}_{\Pi_{n}^{\eta}}\!\left[\boldsymbol{\theta}\right]=\tilde{\boldsymbol{\theta}}_{n}+\frac{1}{\sqrt{s_{n}}}\mathbb{E}\!\left[\boldsymbol{X}_{n}\right],$
so $\sqrt{s_{n}}(\boldsymbol{\theta}_{\mathrm{GB}}-\tilde{\boldsymbol{\theta}}_{n})=\mathbb{E}[\boldsymbol{X}_{n}]$. By Lemma C.1 with $g(\boldsymbol{x})=\boldsymbol{x}$ and the centered Gaussian limit, $\mathbb{E}[\boldsymbol{X}_{n}]\to_{p}\boldsymbol{0}$, proving the first claim.
@@
$s_{n}\operatorname{Var}_{\Pi_{n}^{\eta}}\!\left[\boldsymbol{\theta}\right]=\mathbb{E}\!\left[\boldsymbol{X}_{n}\boldsymbol{X}_{n}^{\top}\right]-\mathbb{E}\!\left[\boldsymbol{X}_{n}\right]\mathbb{E}\!\left[\boldsymbol{X}_{n}\right]^{\top}=\mathbb{E}\!\left[\boldsymbol{X}_{n}\boldsymbol{X}_{n}^{\top}\right]+o_{p}(1),$
since $\mathbb{E}[\boldsymbol{X}_{n}]\to_{p}\boldsymbol{0}$ and $\sup_{n}\mathbb{E}[\boldsymbol{X}_{n}]^{4}<\infty$ by Lemma C.1. ∎
@@
For the asymptotic form of the posterior covariance, fix an admissible center $\tilde{\boldsymbol{\theta}}_{n}$ and consider $\boldsymbol{X}_{n}=\sqrt{s_{n}}(\boldsymbol{\theta}-\tilde{\boldsymbol{\theta}}_{n})$. By Proposition 3.1, the law of $\boldsymbol{X}_{n}$ under $\Pi_{n}^{\eta}$ converges in total variation to $\mathcal{N}(\boldsymbol{0},\boldsymbol{H}_{0}^{-1})$. Lemma C.1 with $g(\boldsymbol{x})=\boldsymbol{x}\boldsymbol{x}^{\top}$ gives
@@
$s_{n}\operatorname{Var}_{\Pi_{n}^{\eta}}\!\left[\boldsymbol{\theta}\right]=\mathbb{E}\!\left[\boldsymbol{X}_{n}\boldsymbol{X}_{n}^{\top}\right]-\mathbb{E}\!\left[\boldsymbol{X}_{n}\right]\mathbb{E}\!\left[\boldsymbol{X}_{n}\right]^{\top}\longrightarrow_{p}\boldsymbol{H}_{0}^{-1},$
@@
$s_{n}\widehat{\boldsymbol{\Sigma}}_{\mathrm{post}}=s_{n}\boldsymbol{\Sigma}_{\mathrm{post},n}+s_{n}\!\left(\widehat{\boldsymbol{\Sigma}}_{\mathrm{post}}-\boldsymbol{\Sigma}_{\mathrm{post},n}\right)\longrightarrow_{p}\boldsymbol{H}_{0}^{-1}.$
@@
By Assumption 2.2 (iii), there exists a neighborhood $\mathfrak{N}$ of $\boldsymbol{\theta}^{\lambda}$ such that $\sup_{\boldsymbol{\theta}\in\mathfrak{N}}\|\boldsymbol{J}_{n}(\boldsymbol{\theta})-\boldsymbol{J}^{\star}\|\to_{p}0$. Since $\bar{\boldsymbol{\theta}}_{n}\to_{p}\boldsymbol{\theta}^{\lambda}$, we have $\boldsymbol{J}_{n}(\bar{\boldsymbol{\theta}}_{n})\to_{p}\boldsymbol{J}^{\star}$. By Assumption 2.1, $\rho$ is $C^{2}$ in a neighborhood of $\boldsymbol{\theta}^{\lambda}$, so $\boldsymbol{H}_{\rho}(\bar{\boldsymbol{\theta}}_{n})\to_{p}\boldsymbol{H}_{\rho}(\boldsymbol{\theta}^{\lambda})$. With $\lambda_{n}\to\lambda$, this yields
@@
Fix an admissible center $\tilde{\boldsymbol{\theta}}_{n}$ and set $\boldsymbol{X}_{n}:=\sqrt{s_{n}}(\boldsymbol{\theta}^{(d)}-\bar{\boldsymbol{\theta}}_{n})$. By Proposition 3.1, conditionally on the data,
$\boldsymbol{X}_{n}\longrightarrow_{d}\mathcal{N}\bigl(\boldsymbol{0},\boldsymbol{H}_{0}^{-1}\bigr),$
@@
Lemma C.2 gives $\sqrt{s_{n}}(\boldsymbol{\theta}_{\mathrm{GB}}-\bar{\boldsymbol{\theta}}_{n})\to_{p}\boldsymbol{0}$, and by the Monte Carlo rate assumption, $\sqrt{s_{n}}(\hat{\boldsymbol{\theta}}_{\mathrm{GB}}-\boldsymbol{\theta}_{\mathrm{GB}})\to_{p}\boldsymbol{0}$. Thus
$\sqrt{s_{n}}\bigl(\boldsymbol{\theta}^{(d)}-\hat{\boldsymbol{\theta}}_{\mathrm{GB}}\bigr)=\boldsymbol{X}_{n}-\sqrt{s_{n}}\bigl(\hat{\boldsymbol{\theta}}_{\mathrm{GB}}-\bar{\boldsymbol{\theta}}_{n}\bigr)=\boldsymbol{X}_{n}+o_{p}(1),$
@@
$\sqrt{s_{n}}\bigl(\boldsymbol{\theta}^{(d)}-\hat{\boldsymbol{\theta}}_{\mathrm{GB}}\bigr)\longrightarrow_{d}\mathcal{N}\bigl(\boldsymbol{0},\boldsymbol{H}_{0}^{-1}\bigr).$
@@
$\sqrt{s_{n}}\bigl(\hat{\boldsymbol{\theta}}_{\mathrm{calib}}^{(d)}-\bar{\boldsymbol{\theta}}_{n}\bigr)=\hat{\boldsymbol{\Omega}}\sqrt{s_{n}}\bigl(\boldsymbol{\theta}^{(d)}-\hat{\boldsymbol{\theta}}_{\mathrm{GB}}\bigr).$
@@
$\sqrt{s_{n}}\bigl(\hat{\boldsymbol{\theta}}_{\mathrm{calib}}^{(d)}-\bar{\boldsymbol{\theta}}_{n}\bigr)\longrightarrow_{d}\mathcal{N}\bigl(\boldsymbol{0},\boldsymbol{\Omega}\boldsymbol{H}_{0}^{-1}\boldsymbol{\Omega}^{\top}\bigr)=\mathcal{N}\bigl(\boldsymbol{0},\boldsymbol{V}_{\mathrm{target}}^{\star}\bigr),$

### v3

@@ Appendix B Proofs of the main results
Write $\varphi(\boldsymbol{x})=N(\boldsymbol{x}\mid\boldsymbol{0},\boldsymbol{H}_{0}^{-1})$. Using the change of variables $\boldsymbol{y}=\boldsymbol{x}+\boldsymbol{\delta}_{n}$ and the triangle inequality,
@@
$\int_{\mathbb{R}^{p}}\Bigl|q_{n}^{\tilde{\boldsymbol{\theta}}_{n}}(\boldsymbol{x})-N\!\bigl(\boldsymbol{x}|\boldsymbol{0},\boldsymbol{H}_{0}^{-1}\bigr)\Bigr|\mathop{}\!\mathrm{d}\boldsymbol{x}\longrightarrow_{p}0.$
Recalling that $\boldsymbol{H}_{0}=\eta\boldsymbol{J}_{\lambda}^{\star}$, this is precisely the claimed total-variation convergence of the law of $s_{n}^{1/2}(\boldsymbol{\theta}-\tilde{\boldsymbol{\theta}}_{n})$ under $\Pi_{n}^{\eta}$ to $N(\boldsymbol{0},(\eta\boldsymbol{J}_{\lambda}^{\star})^{-1})$ at any admissible center $\tilde{\boldsymbol{\theta}}_{n}$. ∎
@@
$s_{n}^{1/2}\bigl(\boldsymbol{\theta}^{(d)}-\boldsymbol{\theta}_{\mathrm{GB}}\bigr)\longrightarrow_{d}N\bigl(\boldsymbol{0},\boldsymbol{H}_{0}^{-1}\bigr),\quad n\to\infty.$ (19)
@@
$s_{n}^{1/2}\bigl(\boldsymbol{\theta}_{\mathrm{calib}}^{(d)}-\tilde{\boldsymbol{\theta}}_{n}\bigr)=\boldsymbol{\Omega}s_{n}^{1/2}\bigl(\boldsymbol{\theta}^{(d)}-\boldsymbol{\theta}_{\mathrm{GB}}\bigr),$ (20)
@@
$s_{n}^{1/2}\bigl(\boldsymbol{\theta}_{\mathrm{calib}}^{(d)}-\tilde{\boldsymbol{\theta}}_{n}\bigr)\longrightarrow_{d}N\bigl(\boldsymbol{0},\boldsymbol{\Omega}\boldsymbol{H}_{0}^{-1}\boldsymbol{\Omega}^{\top}\bigr),\quad n\to\infty.$ (21)
@@
$s_{n}^{1/2}\bigl(\boldsymbol{\theta}_{\mathrm{calib}}^{(d)}-\tilde{\boldsymbol{\theta}}_{n}\bigr)\longrightarrow_{d}N\bigl(\boldsymbol{0},\boldsymbol{V}_{\mathrm{target}}^{\star}\bigr).$
@@ Appendix C Proofs of the plug-in results
Let $q_{n}^{\bar{\boldsymbol{\theta}}_{n}}$ be the density of the scaled posterior $\boldsymbol{X}_{n}=s_{n}^{1/2}(\boldsymbol{\theta}-\tilde{\boldsymbol{\theta}}_{n})$ under $\Pi_{n}^{\eta}$ for an admissible center $\tilde{\boldsymbol{\theta}}_{n}$. We first present two generic facts that will be used repeatedly.
@@
$\int_{\mathbb{R}^{p}}\left|q_{n}^{\bar{\boldsymbol{\theta}}_{n}}(\boldsymbol{x})-\left.N\bigl(\boldsymbol{x}\right|\boldsymbol{0},\boldsymbol{H}_{0}^{-1}\bigr)\right|\mathop{}\!\mathrm{d}\boldsymbol{x}\longrightarrow_{p}0,\quad n\to\infty$
and that there exist random constants $C_{n},c_{n}>0$, bounded in probability, such that, for all $\boldsymbol{x}\in\mathbb{R}^{p}$ and all $n$,
$q_{n}^{\bar{\boldsymbol{\theta}}_{n}}(\boldsymbol{x})\leq C_{n}\exp\!\left(-c_{n}|\boldsymbol{x}|^{2}\right).$
Then, for any function $g:\mathbb{R}^{p}\to\mathbb{R}^{k}$ ($k\in\mathbb{N}$) with polynomial growth,
$\int g(\boldsymbol{x})q_{n}^{\bar{\boldsymbol{\theta}}_{n}}(\boldsymbol{x})\mathop{}\!\mathrm{d}\boldsymbol{x}\longrightarrow_{p}\int g(\boldsymbol{x})\left.N\bigl(\boldsymbol{x}\right|\boldsymbol{0},\boldsymbol{H}_{0}^{-1}\bigr)\mathop{}\!\mathrm{d}\boldsymbol{x}.$
@@
$\mathbb{E}\!\left[\boldsymbol{X}_{n}\right]\longrightarrow_{p}\boldsymbol{0},\quad\mathbb{E}\!\left[\boldsymbol{X}_{n}\boldsymbol{X}_{n}^{\top}\right]\longrightarrow_{p}\boldsymbol{H}_{0}^{-1}.$
@@
On $\{\|\boldsymbol{x}\|\leq R\}$, $g$ is bounded, so total-variation convergence yields convergence of the first term as $n\to\infty$, and the limit as $R\to\infty$ recovers the Gaussian expectation. For the tail term, let
$T_{n,R}:=\int_{\|\boldsymbol{x}\|>R}|g(\boldsymbol{x})|q_{n}^{\bar{\boldsymbol{\theta}}_{n}}(\boldsymbol{x})\mathop{}\!\mathrm{d}\boldsymbol{x}.$
By the Gaussian domination,
$T_{n,R}\leq C_{n}\int_{\|\boldsymbol{x}\|>R}|g(\boldsymbol{x})|\exp(-c_{n}\|\boldsymbol{x}\|^{2})\mathop{}\!\mathrm{d}\boldsymbol{x}.$
Since $g$ has polynomial growth, there exist $m\geq 0$ and $K<\infty$ such that $|g(\boldsymbol{x})|\leq K(1+\|\boldsymbol{x}\|^{m})$. Because $(C_{n})_{n\geq 1}$ and $(c_{n})_{n\geq 1}$ are bounded in probability, for any $\varepsilon>0$ we can choose $B<\infty$ such that
$\sup_{n\geq 1}\operatorname{pr}(C_{n}>B)<\varepsilon,\quad\sup_{n\geq 1}\operatorname{pr}(c_{n}<B^{-1})<\varepsilon.$
On the event $\{C_{n}\leq B,\;c_{n}\geq B^{-1}\}$ we obtain
$T_{n,R}\leq BK\int_{\|\boldsymbol{x}\|>R}(1+\|\boldsymbol{x}\|^{m})\exp\bigl(-B^{-1}\|\boldsymbol{x}\|^{2}\bigr)\mathop{}\!\mathrm{d}\boldsymbol{x}.$
The envelope $(1+\|\boldsymbol{x}\|^{m})\exp(-B^{-1}\|\boldsymbol{x}\|^{2})$ is integrable, so we can choose $R$ large enough that the right-hand side is at most $\varepsilon$. For such $R$,
$\operatorname{pr}(T_{n,R}>\varepsilon)\leq\operatorname{pr}(C_{n}>B)+\operatorname{pr}(c_{n}<B^{-1})\leq 2\varepsilon$
for all $n$, and hence $\sup_{n\geq 1}\Pr(T_{n,R}>\varepsilon)\leq 2\varepsilon$. This shows that $T_{n,R}\to 0$ in probability as $R\to\infty$, uniformly in $n$. This gives convergence of expectations for polynomially growing $g$, and the stated consequences follow by taking $g(\boldsymbol{x})=\boldsymbol{x}$, and $g(\boldsymbol{x})=\boldsymbol{x}\boldsymbol{x}^{\top}$. ∎
@@
$s_{n}^{1/2}\bigl(\boldsymbol{\theta}_{\mathrm{GB}}-\tilde{\boldsymbol{\theta}}_{n}\bigr)\longrightarrow_{p}\boldsymbol{0}$
@@
$s_{n}\operatorname{var}_{\Pi_{n}^{\eta}}\!\left(\boldsymbol{\theta}\right)=\mathbb{E}\!\left[\boldsymbol{X}_{n}\boldsymbol{X}_{n}^{\top}\right]+o_{p}(1),$
where $\boldsymbol{X}_{n}=s_{n}^{1/2}(\boldsymbol{\theta}-\tilde{\boldsymbol{\theta}}_{n})$.
@@
$\boldsymbol{\theta}_{\mathrm{GB}}=\mathbb{E}_{\Pi_{n}^{\eta}}\!\left[\boldsymbol{\theta}\right]=\tilde{\boldsymbol{\theta}}_{n}+s_{n}^{-1/2}\mathbb{E}\!\left[\boldsymbol{X}_{n}\right],$
so $s_{n}^{1/2}(\boldsymbol{\theta}_{\mathrm{GB}}-\tilde{\boldsymbol{\theta}}_{n})=E(\boldsymbol{X}_{n})$. By Lemma C.1 with $g(\boldsymbol{x})=\boldsymbol{x}$ and the centered Gaussian limit, $E(\boldsymbol{X}_{n})\to_{p}\boldsymbol{0}$, proving the first claim.
@@
$s_{n}\operatorname{var}_{\Pi_{n}^{\eta}}\!\left(\boldsymbol{\theta}\right)=\mathbb{E}\!\left[\boldsymbol{X}_{n}\boldsymbol{X}_{n}^{\top}\right]-\mathbb{E}\!\left[\boldsymbol{X}_{n}\right]\mathbb{E}\!\left[\boldsymbol{X}_{n}\right]^{\top}=\mathbb{E}\!\left[\boldsymbol{X}_{n}\boldsymbol{X}_{n}^{\top}\right]+o_{p}(1),$
since $E(\boldsymbol{X}_{n})\to_{p}\boldsymbol{0}$ by Lemma C.1. ∎
@@
For the asymptotic form of the posterior covariance, fix an admissible center $\tilde{\boldsymbol{\theta}}_{n}$ and consider $\boldsymbol{X}_{n}=s_{n}^{1/2}(\boldsymbol{\theta}-\tilde{\boldsymbol{\theta}}_{n})$. By Proposition 3.1, the law of $\boldsymbol{X}_{n}$ under $\Pi_{n}^{\eta}$ converges in total variation to $N(\boldsymbol{0},\boldsymbol{H}_{0}^{-1})$. Lemma C.1 with $g(\boldsymbol{x})=\boldsymbol{x}\boldsymbol{x}^{\top}$ gives
@@
$s_{n}\operatorname{var}_{\Pi_{n}^{\eta}}\!\left(\boldsymbol{\theta}\right)=\mathbb{E}\!\left[\boldsymbol{X}_{n}\boldsymbol{X}_{n}^{\top}\right]-\mathbb{E}\!\left[\boldsymbol{X}_{n}\right]\mathbb{E}\!\left[\boldsymbol{X}_{n}\right]^{\top}\longrightarrow_{p}\boldsymbol{H}_{0}^{-1},$
@@
$s_{n}\hat{\boldsymbol{\Sigma}}_{\mathrm{post}}=s_{n}\boldsymbol{\Sigma}_{\mathrm{post},n}+s_{n}\!\left(\hat{\boldsymbol{\Sigma}}_{\mathrm{post}}-\boldsymbol{\Sigma}_{\mathrm{post},n}\right)\longrightarrow_{p}\boldsymbol{H}_{0}^{-1}.$
@@
By Assumption 2.2 (iii), there exists a neighborhood $\mathcal{N}$ of $\boldsymbol{\theta}^{\lambda}$ such that $\sup_{\boldsymbol{\theta}\in\mathcal{N}}\|\boldsymbol{J}_{n}(\boldsymbol{\theta})-\boldsymbol{J}^{\star}\|\to_{p}0$. Since $\bar{\boldsymbol{\theta}}_{n}\to_{p}\boldsymbol{\theta}^{\lambda}$, we have $\boldsymbol{J}_{n}(\bar{\boldsymbol{\theta}}_{n})\to_{p}\boldsymbol{J}^{\star}$. By Assumption 2.1, $\rho$ is $C^{2}$ in a neighborhood of $\boldsymbol{\theta}^{\lambda}$, so $\boldsymbol{H}_{\rho}(\bar{\boldsymbol{\theta}}_{n})\to_{p}\boldsymbol{H}_{\rho}(\boldsymbol{\theta}^{\lambda})$. With $\lambda_{n}\to\lambda$, this yields
@@
Fix an admissible center $\tilde{\boldsymbol{\theta}}_{n}$ and set $\boldsymbol{X}_{n}:=s_{n}^{1/2}(\boldsymbol{\theta}^{(d)}-\bar{\boldsymbol{\theta}}_{n})$. By Proposition 3.1, conditionally on the data,
$\boldsymbol{X}_{n}\longrightarrow_{d}N\bigl(\boldsymbol{0},\boldsymbol{H}_{0}^{-1}\bigr),$
@@
Lemma C.2 gives $s_{n}^{1/2}(\boldsymbol{\theta}_{\mathrm{GB}}-\bar{\boldsymbol{\theta}}_{n})\to_{p}\boldsymbol{0}$, and by the Monte Carlo rate assumption, $s_{n}^{1/2}(\hat{\boldsymbol{\theta}}_{\mathrm{GB}}-\boldsymbol{\theta}_{\mathrm{GB}})\to_{p}\boldsymbol{0}$. Thus
$s_{n}^{1/2}\bigl(\boldsymbol{\theta}^{(d)}-\hat{\boldsymbol{\theta}}_{\mathrm{GB}}\bigr)=\boldsymbol{X}_{n}-s_{n}^{1/2}\bigl(\hat{\boldsymbol{\theta}}_{\mathrm{GB}}-\bar{\boldsymbol{\theta}}_{n}\bigr)=\boldsymbol{X}_{n}+o_{p}(1),$
@@
$s_{n}^{1/2}\bigl(\boldsymbol{\theta}^{(d)}-\hat{\boldsymbol{\theta}}_{\mathrm{GB}}\bigr)\longrightarrow_{d}N\bigl(\boldsymbol{0},\boldsymbol{H}_{0}^{-1}\bigr).$
@@
$s_{n}^{1/2}\bigl(\hat{\boldsymbol{\theta}}_{\mathrm{calib}}^{(d)}-\bar{\boldsymbol{\theta}}_{n}\bigr)=\hat{\boldsymbol{\Omega}}s_{n}^{1/2}\bigl(\boldsymbol{\theta}^{(d)}-\hat{\boldsymbol{\theta}}_{\mathrm{GB}}\bigr).$
@@
$s_{n}^{1/2}\bigl(\hat{\boldsymbol{\theta}}_{\mathrm{calib}}^{(d)}-\bar{\boldsymbol{\theta}}_{n}\bigr)\longrightarrow_{d}N\bigl(\boldsymbol{0},\boldsymbol{\Omega}\boldsymbol{H}_{0}^{-1}\boldsymbol{\Omega}^{\top}\bigr)=N\bigl(\boldsymbol{0},\boldsymbol{V}_{\mathrm{target}}^{\star}\bigr),$